# FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with Multimodal Large Language Models

### v1

@@ 5.2 Implementation Details
5.3 Comparison with Existing Methods
5.3.1 Coarse-Grained Classification Task
For the coarse-grained classification task, we compare FaceShield with state-of-the-art FAS methods  [19, 59, 77, 78] and open-source MLLM models [38, 4, 79, 20].
Intra-dataset Testing. Table 1 demonstrates that Our FaceShield significantly outperforms three representative traditional FAS methods [19, 59, 77] in within-domain experiments. Moreover, our performance greatly exceeds the zero-shot capabilities of general MLLM. We also fine-tune the open-source MLLM (i.e., Bunny [20]), selecting RGB images and language data from the dataset to conduct experiments on Bunny. We find that FaceShield also surpasses the well-tuned MLLM (Bunny) with 1% HTER decrease.
Cross-dataset Testing. Table 2 shows the performance of FaceShield in cross-domain scenarios, where we trained on two out of three selected datasets and tested on one. FaceShield demonstrates performance far exceeding traditional FAS models in cross-domain scenarios. Under the S&PW protocol, it achieves the HTER of 5.72%, showcasing FaceShield’s strong generalization capabilities compared to traditional methods.
5.3.2 Fine-Grained Classification Task
Table 3 shows the results under the fine-grained classification task. For open-source MLLMs [38, 4, 79, 20], we incorporated 12 types of attacks into the prompt, allowing it to respond with the correct type. For the fine-tuned MLLM and our FaceShield, we selected keywords from the responses for evaluation. It is evident that supervised fine-tuning can significantly improve the model’s performance, with FaceShieldachieving the best results. Cross-dataset results can be found in the Appendix due to space constraints.
5.3.3 Reasoning Task
5.3.4 Attack Localization Task
@@
Qwen-VL [4] 2.07 1.49
Lenna [63] 37.77 35.41
Sphinx [35] 47.86 46.30
Bunny [20] 73.50 71.65
Bunny(fine-tuned) [20] 92.30 89.71
@@
$\checkmark$ $\checkmark$ $\mathbf{99.41\pm 0.06}$ $\mathbf{0.53\pm 0.06}$ $\mathbf{95.81\pm 0.11}$ $99.29\pm 0.04$ $0.57\pm 0.04$ $\mathbf{97.78\pm 0.21}$ $\mathbf{95.6\pm 0.19}$
Figure 6: Comparison of performance after fine-tuning using our proposed dataset on LLaVA [38] and Bunny [20] models.
5.4 Ablation Study
Effectiveness of the proposed datasets. We conduct pre-training and SFT with our proposed FaceShield-pre10K and FaceShield-sft45K on LLaVA [38] and Bunny [20] to evaluate the efficacy and generalization of our instruction tuning datasets. The results in Fig. 6 demonstrate that fine-tuning on our dataset significantly enhances performance across three tasks for both LLaVA and Bunny. This demonstrates that our dataset can effectively enrich general MLLMs with substantial FAS-related knowledge and boost their capabilities in FAS tasks. It also validates the efficacy of our advanced dataset construction pipeline.
Figure 7: Ablation of visual token masking ratio $p$ in PVTM on (a) three tasks (i.e., coarse- & fine-grained classification, and reasoning), and (b) attack localization task.
@@ 5.4 Ablation Study
Figure 8: Ablation of visual token retain ratio $k$ in PVTM on (a) three tasks (i.e., coarse- & fine-grained classification, and reasoning), and (b) attack localization task.
@@
We further conduct studies of PVTM with different visual token masking ratios $p$, as shown in Fig. 7(a) and (b). Through our experiments for three (coarse- and fine-grained classification, and attack localization) tasks, we find that masking 5% of the tokens strikes an optimal balance between reducing spurious correlations and preserving essential information. However, we cannot find any improvement when masking more tokens due to severe information loss. Besides, we also investigate varying proportions of visual token retain ratio $k$, as illustrated in Fig. 8(a) and (b). We preserve 0%, 10%, 20%, and 30% of the tokens, respectively, and then randomly mask $p$=5the remaining tokens. The results show that retaining $k$=10visual tokens with strong importance achieves optimal performance. However, the more tokens are preserved, the poorer the model performs, suggesting that as the importance of tokens decreases, the likelihood of spurious correlations increases.
Figure 9: Importance visualization of (a) RGB vision tokens for classification tasks and (b) SAV tokens for attack localization task.
6 Visualization and Analysis
Here, we visualize the importance of RGB and SAV tokens. Fig. 9(a) displays the significance of visual tokens in relation to the prompt text for images from three different categories. For the coarse- and fine-grained classification tasks, it is evident that the tokens sensitive to prompt are mainly clustered in the facial regions. The features of these areas play a decisive role in determining the final outcome. Fig. 9(b) demonstrates how SAV tokens assist in the attack localization task. For this task, RGB tokens sometimes fail to provide accurate localization information. As shown in the left of Fig. 9(b), for the eye mask region, the importance visualization of RGB tokens is quite dispersed. However, in the SAV token visualization, the significance of the eye attack region is clearly highlighted. The right part of Fig. 9(b) shows how RGB and SAV tokens together provide supportive information for the attack localization task, including the eye mask attack area and the holding hands.
7 Conclusion

### v2

@@ Introduction
Qwen-VL (Qwen-VL) 2.07 1.49
Lenna (wei2023lenna) 37.77 35.41
Sphinx (lin2023sphinxjointmixingweights) 47.86 46.30
Bunny (he2024bunny) 73.50 71.65
Bunny(fine-tuned) (he2024bunny) 92.30 89.71
@@
Table 8: Ablation study on pretraining w/ or w/o FaceShield-pre10K dataset.
FaceShield-pre10K Fine-grained Classification Reasoning
ACC (%) $\uparrow$ ACC (%) $\uparrow$ HTER (%) $\downarrow$
$\times$ 94.78 $\pm$ 0.19 98.83 $\pm$ 0.06 0.94 $\pm$ 0.05
$\checkmark$ 95.81 $\pm$ 0.11 99.29 $\pm$ 0.04 0.57 $\pm$ 0.04
Intra-dataset Testing.
Table 2 demonstrates that Our FaceShield significantly outperforms three representative traditional FAS methods (he2016deep; Wang_2022_CVPR; zhou2022learning). Moreover, our performance greatly exceeds the zero-shot capabilities of general MLLM. We also fine-tune the open-source MLLM (i.e., Bunny (he2024bunny)), selecting RGB images and language data from the dataset to conduct experiments on Bunny. We find that FaceShield also surpasses the well-tuned MLLM (Bunny) with 1% HTER decrease.
Cross-dataset Testing.
Table 3 shows the performance of FaceShield in cross-domain scenarios, where we trained on two out of three selected datasets and tested on one. FaceShield demonstrates performance far exceeding traditional FAS models in cross-domain scenarios. Under the S&P→W protocol, it achieves the HTER of 5.72%, showcasing FaceShield’s strong generalization capabilities compared to traditional methods. Further results in Table 4 confirm its robustness across unseen datasets like CASIA-MFSD and Replay-Attack.
@@
$\checkmark$ $\checkmark$ 99.41 $\pm$ 0.06 0.53 $\pm$ 0.06 95.81 $\pm$ 0.11 99.29 $\pm$ 0.04 0.57 $\pm$ 0.04 97.78 $\pm$ 0.21 95.6 $\pm$ 0.19
Fine-Grained Classification Task
Table 5 shows the results under the fine-grained classification task. For open-source MLLMs, we incorporated 12 types of attacks into the prompt, allowing it to respond with the correct type. For the fine-tuned MLLM and our FaceShield, we selected keywords from the responses for evaluation. It is evident that supervised fine-tuning can significantly improve the model’s performance, with FaceShield achieving the best results.
Reasoning Task
Attack Localization Task
Figure 5: Comparison of performance after fine-tuning using our proposed dataset on LLaVA and Bunny models.
Figure 6: Ablation of visual token masking ratio $p$ in PVTM on three tasks (i.e., coarse- & fine-grained classification, and reasoning)
Figure 7: Ablation of visual token retain ratio $k$ in PVTM on three tasks (i.e., coarse- & fine-grained classification, and reasoning)
Ablation Study
Effectiveness of the proposed datasets. We conduct pre-training and SFT with our proposed FaceShield-pre10K and FaceShield-sft45K on LLaVA (liu2023llava) and Bunny (he2024bunny) to evaluate the efficacy and generalization of our constructed datasets. As shown in Table 8, pretraining with the FaceShield-pre10K dataset significantly improves performance, with fine-grained classification accuracy and reasoning accuracy increasing, while HTER is reduced. This demonstrates that pretraining with FaceShield-pre10K enhances the model’s capabilities in FAS-related tasks.
Additionally, the results in Fig. 5 show that fine-tuning on our dataset further boosts performance across three tasks for both LLaVA and Bunny. This validates the effectiveness of our dataset in enriching MLLMs with FAS-related knowledge and improving their overall performance in FAS tasks, supporting the efficacy of our advanced dataset construction pipeline.
@@
We further study PVTM with different visual token masking ratios $p$, as shown in Fig.6. Through experiments on three (coarse- and fine-grained classification, and reasoning) tasks, we find that masking 5% of the tokens strikes an optimal balance between reducing spurious correlations and preserving essential information. However, no improvement is found when masking more tokens due to severe information loss. Additionally, we investigate varying proportions of visual token retain ratio $k$, as illustrated in Fig.7. We preserve 0%, 10%, 20%, and 30% of the tokens, respectively, and then randomly mask pp=5% of the remaining tokens. The results show that retaining kk=10% visual tokens with strong importance achieves optimal performance. However, the more tokens preserved, the poorer the model performs, suggesting that as the importance of tokens decreases, the likelihood of spurious correlations increases.
Visualization and Analysis
Figure 8: Importance visualization of SAV tokens for attack localization task.
Fig. 8 illustrates the visualization obtained after applying SAVP to the attack localization task, highlighting the effectiveness of SAV tokens. Compared to RGB tokens, which may suffer from dispersed attention, SAV tokens, once applied, can accurately focus on spoofed regions (e.g., eyeglass areas, hand-held masks). These tokens show strong alignment with the ground-truth annotations, producing more focused and interpretable activations around spoof artifacts. Additionally, visual tokens in the attack regions exhibit higher similarity scores with the task prompt. With PVTM applied, the model is further guided to focus on these deception-relevant areas, emphasizing the critical roles of both SAV tokens and PVTM in improving localization performance.
Conclusion
@@
FaceShield[Uncaptioned image]: Explainable Face Anti-Spoofing with
Multimodal Large Language Models
Appendix A Details about FaccShield Dataset
Data Source
The FaceShield dataset, comprising FaceShield-pre10K for pretraining and FaceShield-sft45K for supervised fine-tuning (SFT), is constructed using images sourced from three widely recognized datasets: WMCA (8714076), SiW-Mv2 (xiaoguo2022MDFAS), and PADISI (Rostami_2021_ICCV). These foundational datasets provide a diverse basis for generating question-answer (QA) pairs that address a variety of tasks in anti-spoofing research.
In the pretraining phase, the FaceShield-pre10K dataset focuses on generating QA pairs that describe the visual content of the images. A total of 9,297 QA pairs were created to establish a foundational understanding of the visual data. These pairs enable models to grasp core visual concepts and perform basic reasoning related to spoofing scenarios, forming the groundwork for subsequent learning.
In contrast, the FaceShield-sft45K dataset, used in the SFT phase, incorporates task-specific annotations to address more complex objectives. This dataset contains 45,662 QA pairs, which are categorized into several distinct tasks: coarse-grained classification, fine-grained classification, reasoning, and attack localization. These QA pairs are designed to meet practical anti-spoofing requirements, enabling models to handle diverse challenges across various anti-spoofing tasks.
The FaceShield-pre10K dataset emphasizes descriptive comprehension and fundamental reasoning about image content, while the FaceShield-sft45K dataset advances to specialized, task-driven annotations. The distribution of QA pairs is detailed in Table 10 and the category distributions in Fig. 9.
Table. 11 compares face anti-spoofing datasets. Unlike traditional datasets with only binary labels for print and replay attacks, FaceShield provides richer annotations (including attack types and bounding boxes) and supports multimodal tasks with large-scale image-caption and QA pairs.
Figure 9: Comparison of category distributions in FaceShield-pre10K and FaceShield-sft45K datasets.
Table 10: Dataset Statistics(QA pairs)
Datasets Source Data Size Total
WMCA (8714076) 3875
FaceShield-pre10K SiW-Mv2 (xiaoguo2022MDFAS) 3782 9297
PADISI (Rostami_2021_ICCV) 1640
WMCA (8714076) 16776
FaceShield-sft45K SiW-Mv2 (xiaoguo2022MDFAS) 18096 45662
PADISI (Rostami_2021_ICCV) 10790
Table 11: Comparison of existing datasets
Dataset Attack Types Annotations
SiW Print, Replay(2) Binary class
Oulu-NPU Print, Replay(2) Binary class
CASIA-MFSD Print, Replay(2) Binary class
Replay-Attack Print, Replay(2) Binary class
MSU Print, Replay(2) Binary class
ROSE Print, Replay, PaperMask(3) Binary class
FaceShield Unified-attack (11 types) QA and bounding boxes
Data Format
We constructed the dataset based on the LLAVA framework, as shown in Fig. 10.
Figure 10: Data format
Details about FaceShield-pre10K
Data Generation
As shown in Fig. 11, we employed Bunny-Llama-3-8B-V (MLLM) (he2024bunny) to generate high-quality question-answer (QA) pairs aimed at describing the visual content of images. The primary objective of this dataset is to equip the MLLM with a foundational understanding of visual attributes relevant to the Face Anti-Spoofing (FAS) task, enabling it to develop basic perceptual capabilities to distinguish between real and spoofed faces.
Figure 11: FaceShield-pre10Kconstruction pipeline
The QA generation process is guided by a systematically designed prompt framework comprising three key components. First, each image is paired with its corresponding Ground Truth (GT), which denotes the spoofing category (e.g., glasses, mask, makeup). The GT serves as contextual information to ensure that the generated QA pairs are semantically aligned with the spoofing type depicted in the image. Second, the construction requirements are explicitly defined to standardize the format and content of the QA pairs. Each question is concise, typically a single sentence, and designed to elicit descriptions focused on either the environment (e.g., background, lighting) or facial attributes (e.g., gender, facial organ, expression) (eccv25tffas). The answer, by contrast, is required to be detailed and comprehensive, integrating observations about both the environment and facial features. This approach ensures that the QA pairs are both informative and task-relevant. Lastly, the prompts include illustrative examples to provide clear guidance on the desired output structure and level of detail. These examples demonstrate the expected question format, the richness required in the answers, and the integration of diverse attributes such as lighting conditions and facial expressions. This structured approach ensures that the generated QA pairs are of high quality, semantically relevant, and tailored to the needs of the FAS domain. Consequently, the FaceShield-pre10K dataset provides a robust foundation for the MLLM to develop perceptual capabilities, facilitating its adaptation to more complex and specialized tasks in the subsequent fine-tuning phase.
Data Filtering
To ensure the quality of the generated QA pairs, we utilized the CLIP (clip) model to evaluate the semantic alignment between the images and their corresponding textual descriptions. The filtering process was based on similarity scores computed between each image and its associated text. Samples with a similarity score below 15% were considered low-quality and excluded from the dataset, while the rest were retained.
For text data exceeding the 77-token limit of CLIP (clip), the text was segmented into multiple sentences, and similarity scores were computed for each sentence. The final similarity score was calculated as the average of the sentence-level scores. For text samples within the token limit, the similarity score was directly computed using the entire text. This approach ensured accurate evaluation for longer text samples while maintaining consistency for shorter ones.
After applying this filtering process, the dataset was reduced from 12091 to 9297 QA pairs. This step ensured that the dataset retained only high-quality, semantically consistent data, forming a reliable basis for pretraining tasks in the FAS domain.
Details about FaceShield-sft45K
As shown in Fig. 12, we utilized Bunny-Llama-3-8B-V (he2024bunny) to generate high-quality QA pairs for four tasks: Coarse-grained Classification, Fine-grained Classification, Reasoning, and Attack Localization. The prompts for data generation follow a common structure comprising Ground Truth (GT), Requirements, and Examples. GT provides contextual information (e.g., real or spoof labels, spoof types, and bounding box coordinates), Requirements specify the desired QA format, and Examples guide generation with representative samples. While the structure is consistent, the content is tailored to the objectives of each task. Data sizes for each task are detailed in Table 12.
Figure 12: FaceShield-sft45K construction pipeline
The construction process, outlined in Algorithm 1, includes generating QA pairs using multimodal input, quality filtering, human review, and data augmentation through rephrasing. This pipeline ensures a diverse, high-quality dataset optimized for the specified tasks.
Table 12: Task and Data Size(QA pairs) for FaceShield-sft45K Dataset
Dataset Task Data Size
FaceShield-sft45K Coarse-grained Classification 13076
Fine-grained Classification 12749
Reasoning 11082
Attack Localization 8755
Multi-task Data generation
Coarse-grained Classification Task
The coarse-grained classification task aims to determine whether a face in an image is real or spoofed. This task addresses the fundamental binary classification problem central to FAS. In this task, the prompts focus on eliciting clear and definitive judgments about the authenticity of the face, supported by reasoning tied to visual features. For example, like Fig. 13, a typical prompt requires the MLLM to generate a question such as, “Is the face in the picture real or spoof?” and an answer like, “The face in the picture is a spoof face. This is because the face appears glossy, has an unusual facial reflex, and therefore it is part of an image played on an electronic device.” This task provides the model with foundational knowledge to identify face authenticity.
Fine-grained Classification Task
The fine-grained classification task extends the binary classification by requiring the model to identify not only whether the face is real or spoofed but also the specific spoof type (e.g., glasses, fakehead, makeup attack). The prompts for this task are designed to ensure the generated QA pairs explicitly address these finer distinctions. For example, like Fig. 14, a question might ask, “What attack type of facial spoof is shown in this image?” with an answer such as, “The image showcases a Glasses Attack, where the subject’s face is cleverly obscured by a pair of glasses.” This task enhances the model’s ability to detect and classify diverse spoofing techniques.
Reasoning Task
The reasoning task is designed to enhance the model’s ability to provide structured and interpretable explanations for its judgments, fostering a deeper understanding of FAS scenarios. For instance, like Fig. 15, the question might be phrased as: “Why is the face real or spoofed?” The answer requires the model to systematically evaluate the image based on four key attributes: Facial Lighting, Global Shape Consistency of Facial Features, Sense of Depth and Three-Dimensionality, and Presence of Phone Screen or Paper Edges. These attributes are selected because they represent core indicators in FAS tasks. Facial Lighting highlights irregularities caused by artificial materials or spoofing media under light. Global Shape Consistency assesses whether the facial features align naturally, a key distinction often disrupted in spoofing attempts. Sense of Depth and Three-Dimensionality captures the lack of realistic depth commonly found in spoofed faces. Presence of Phone Screen or Paper Edges identifies visible edges or surfaces indicative of physical spoofing mediums, such as photographs or displays. By methodically analyzing each of these attributes and integrating their observations, the model delivers a comprehensive judgment on the authenticity of the face.
Attack Localization Task
The attack localization task focuses on identifying the specific region of a spoofing attack within an image, which is critical for detailed analysis and interpretability in FAS scenarios. The prompts require the model to determine whether a face is real or spoofed and, if spoofed, to locate the spoofing region. For instance, like Fig. 16, the generated QA might include a question like, “Can you locate the spoof area of the face in the image?” The corresponding answer should provide the bounding box (bbox) coordinates of the spoofing region (e.g., eyes, mouth) and a concise description of the observed spoofing characteristics. This task ensures that the model not only identifies the presence of a spoof but also pinpoints its location, thereby improving interpretability in FAS decision-making.
Figure 13: Prompt for Coarse-grained Classification Task
Figure 14: Prompt for Fine-grained Classification Task
Figure 15: Prompt for Reasoning Task
Figure 16: Prompt for Attack Localization Task
Data Filtering
To ensure the quality and task relevance of the QA pairs in the FaceShield-sft45K dataset, a multi-stage filtering process was employed, incorporating LLM-based evaluation, special words filtering, and human review. This pipeline ensured that low-quality or irrelevant QA pairs were systematically identified and removed, resulting in a high-quality dataset suitable for supervised fine-tuning in Face Anti-Spoofing (FAS) tasks.
LLM-based Filtering
The first stage of filtering utilizes an LLM to evaluate the semantic consistency between each question and its corresponding answer. The LLM was tasked with determining whether the answer directly addressed the question and adhered to the specific requirements of the task. QA pairs were flagged for exclusion if:
  1. 1.
    The question and answer exhibited semantic misalignment or inconsistency.
  2. 2.
    The answer deviated from the task objective or provided irrelevant information.
This automated evaluation allowed us to efficiently filter out QA pairs that lacked coherence or failed to meet the standards of task relevance.
Special Words Filtering
In parallel, a second filtering mechanism was implemented to identify QA pairs containing predefined special words or phrases indicative of low-quality responses. These phrases typically signaled either the model’s inability to provide a meaningful answer or responses irrelevant to the task. Examples of such phrases include:
  • “As an AI language model, I cannot…”
  • “It is difficult to determine…”
  • “It is not possible to…”
QA pairs containing such phrases were automatically removed, as they failed to provide informative or useful annotations for the dataset.
Human Review and Refinement
Following the automated filtering stages, human reviewers conducted a final review of the retained QA pairs to ensure quality and consistency. This review process focused on:
  1. 1.
    Addressing cases where automated filtering may have incorrectly flagged or missed low-quality QA pairs.
  2. 2.
    Correcting minor errors, such as grammatical inconsistencies or subtle misalignments between questions and answers.
  3. 3.
    Refining edge cases that required nuanced contextual understanding.
The human review provided an additional layer of quality assurance, ensuring that the dataset met the required standards of accuracy, relevance, and clarity.
Data Augmentation
To further increase the diversity and quantity of QA pairs in the FaceShield-sft45K dataset, a systematic data augmentation strategy that leverages the generative capabilities of the LLM was employed. As shown in Fig. 17, this method used the existing QA pairs as input and generated additional semantically equivalent or closely related QA pairs, enhancing linguistic variability while maintaining the original semantic intent.
The augmentation process was structured as follows: the LLM was provided with a QA pair in a structured format, including both the question and answer. The prompt instructed the LLM to act as an expert in data augmentation, generating multiple new QA pairs that conveyed the same meaning as the original but differed in phrasing, word choice, and structural elements. The generated outputs were designed to reflect diverse linguistic expressions while ensuring semantic alignment with the original QA pair.
By applying this augmentation process consistently across the dataset, the FaceShield-sft45K dataset was significantly enriched with variations that introduced natural linguistic diversity. This augmentation ensured that the dataset could support robust model training by exposing the models to a wider range of expressions and syntactic patterns, improving their ability to generalize across varied language inputs.
Figure 17: Prompt for Data augmentation
Ablation Study
This section investigates the contributions of key components in the training pipeline, specifically the inclusion of the FaceShield-pre10K dataset during pretraining and the data filtering and augmentation processes during fine-tuning. The ablation experiments provide quantitative insights into the significance of these components.
To evaluate the impact of pretraining on the FaceShield-pre10K dataset, we compared the performance of FaceShield pretrained with and without this dataset. As shown in Table 13, pretraining with FaceShield-pre10K improved fine-grained classification accuracy from $94.78\%\pm 0.19$ to $95.81\%\pm 0.11$, and reasoning accuracy from $98.83\%\pm 0.06$ to $99.29\%\pm 0.04$. Moreover, the Half Total Error Rate (HTER) was significantly reduced from $0.94\%\pm 0.05$ to $0.57\%\pm 0.04$, highlighting the dataset’s effectiveness in improving both classification and interpretative capabilities.
The influence of data filtering and augmentation during supervised fine-tuning was analyzed using three configurations, as summarized in Table 14. Without any filtering or augmentation, the model achieved a fine-grained classification accuracy of $87.92\%\pm 0.16$, reasoning accuracy of $98.14\%\pm 0.06$, and an HTER of $1.63\%\pm 0.05$. Adding data filtering improved these metrics to $90.88\%\pm 0.19$, $98.73\%\pm 0.03$, and $1.24\%\pm 0.08$, respectively. When both filtering and augmentation were applied, the model achieved the best performance with fine-grained classification accuracy of $95.81\%\pm 0.11$, reasoning accuracy of $99.29\%\pm 0.04$, and an HTER reduced to $0.57\%\pm 0.04$. These results underscore the complementary effects of filtering, which ensures task-relevant and semantically consistent data, and augmentation, which introduces linguistic variability and improves generalization.
Table 13: Ablation Study on Pretraining w/ or w/o FaceShield-pre10K Dataset.
FaceShield-pre10K Fine-grained Classification Reasoning
ACC (%) $\uparrow$ ACC (%) $\uparrow$ HTER (%) $\downarrow$
$\times$ 94.78 $\pm$ 0.19 98.83 $\pm$ 0.06 0.94 $\pm$ 0.05
95.81 $\pm$ 0.11 99.29 $\pm$ 0.04 0.57 $\pm$ 0.04
Table 14: Ablation Study on Data filtering and Data augmentation.
Data filtering Data augmentation Fine-grained classification Reasoning
ACC(%) $\uparrow$ ACC(%) $\uparrow$ HTER(%) $\downarrow$
$\times$ $\times$ 87.92 $\pm$ 0.16 98.14 $\pm$ 0.06 1.63 $\pm$ 0.05
$\times$ 90.88 $\pm$ 0.19 98.73 $\pm$ 0.03 1.24 $\pm$ 0.08
95.81 $\pm$ 0.11 99.29 $\pm$ 0.04 0.57 $\pm$ 0.04
Overall, the ablation study demonstrates the critical role of pretraining on FaceShield-pre10K, which provides foundational knowledge for downstream tasks, and the combined use of data filtering and augmentation, which enhances both data quality and linguistic diversity. By isolating these components, we confirm the robustness and effectiveness of the proposed training pipeline for multimodal anti-spoofing tasks.
Appendix B Evaluation Metrics
To comprehensively evaluate the performance of the proposed model across different tasks, this study employs a range of evaluation metrics, including Half Total Error Rate (HTER) (yu2022deep), Accuracy (ACC), BLEU (papineni2002bleu), ROUGE-L (lin2004rouge), METEOR (banerjee2005meteor), and Average Precision (AP@40 and AP@50).
The choice of these evaluation metrics is motivated by the specific nature of each task. ACC and HTER are commonly used in Face Anti-Spoofing (FAS) tasks, providing a direct measure of model performance in distinguishing between genuine and spoofed faces. For the reasoning task, which involves generating long textual content for analysis, metrics such as BLEU, ROUGE-L, and METEOR are suitable as they are widely used in Natural Language Processing (NLP) to evaluate the quality of generated text. Lastly, for the localization task, where the goal is to locate the spoof area, we include object detection metrics like AP@40 and AP@50, which are effective in measuring the accuracy of the predicted regions of interest in detecting spoofed areas.
The detailed definitions and calculations of these metrics are presented as follows:
Half Total Error Rate (HTER)
HTER is a critical metric for binary classification tasks, especially in scenarios with imbalanced data. It is defined as:
$\text{HTER}=\frac{\text{FAR}+\text{FRR}}{2},$ (9)
where FAR (False Acceptance Rate) represents the proportion of negative samples that are incorrectly classified as positive, and FRR (False Rejection Rate) represents the proportion of positive samples that are incorrectly classified as negative.
Accuracy (ACC)
Accuracy evaluates the proportion of correctly classified samples among all samples. It is defined as:
$\text{ACC}=\frac{\text{TP}+\text{TN}}{\text{Total}},$ (10)
where TP and TN are the true positives and true negatives, respectively, and Total is the sum of all samples.
BLEU (Bilingual Evaluation Understudy)
BLEU is a metric commonly used in natural language generation tasks to evaluate the n-gram overlap between generated and reference texts. In this study, we utilize BLEU-1, BLEU-2, BLEU-3, and BLEU-4, which correspond to the 1-gram, 2-gram, 3-gram, and 4-gram match rates, respectively. The general formula for BLEU is:
$\text{BLEU-N}=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$ (11)
where $N$ is the maximum n-gram length, $p_{n}$ is the proportion of matched n-grams, $w_{n}$ is the weight (usually equally distributed), and BP is the brevity penalty to adjust for short generated texts:
$\text{BP}=\begin{cases}1&\text{if }c>r,\\ e^{(1-r/c)}&\text{if }c\leq r,\end{cases}$ (12)
where $c$ and $r$ are the lengths of the generated text and the reference text, respectively.
ROUGE-L (Recall-Oriented Understudy for Gisting Evaluation)
ROUGE-L is based on the longest common subsequence (LCS) and measures the similarity between the generated and reference texts in terms of sequential word matches. It is defined as:
$\text{ROUGE-L}=\frac{\text{LCS}(\text{gen},\text{ref})}{\max(\text{len}(\text{gen}),\text{len}(\text{ref}))},$ (13)
where LCS is the length of the longest common subsequence between the generated text gen and the reference text ref, and $\text{len}(\cdot)$ denotes the length of the respective text.
METEOR (Metric for Evaluation of Translation with Explicit ORdering)
METEOR evaluates the semantic similarity between the generated and reference texts by considering word matching, stemming, and synonymy. It is computed as:
$\text{METEOR}=F_{\text{mean}}\cdot(1-\text{Penalty}),$ (14)
where $F_{\text{mean}}$ is the harmonic mean of precision and recall, and Penalty is a factor penalizing repeated patterns.
Appendix C Memory Requirement
As shown in Tab. 15, the FaceShield model, with 3.83B parameters and a size of 10.42 GB, allocates 7.14 GB of memory during inference. This model achieves a balance between parameter count, model size, and memory allocation, making it suitable for environments with limited resources while maintaining strong performance.
Model Parameters Model Size Memory Allocated
FaceShield 3.83B 10.42 GB 7.14 GB
Table 15: Model Specifications
Appendix D Detailed Experiments Setting
Detailed Setting of Model Architectures
As illustrated in Fig. 4(a) depicts the baseline model, similar to LLaVA, comprising key components such as the Vision Encoder, Projector, Tokenizer, and LLM. This model only accepts RGB modality images as input. However, the high visual similarity between genuine faces and presentation attacks (PAs) in the RGB domain presents a significant challenge for this approach. In Fig. 4(b), the input is extended to include auxiliary modality images, which corresponds to the proposed SAVP strategy. Specifically, the auxiliary modality images are generated by applying operations such as Gray, HOG, and LBP to the RGB images, thereby enhancing the model’s ability to perceive attack cues. In Fig. 4(c), we introduce the final model, which incorporates the PVTM module. This module retains and masks visual tokens based on the similarity between visual tokens and textual tokens, thereby reducing interference and redundancy. For a detailed explanation of the PVTM module, refer to Fig. 4(c).
Overall, the framework consists of a Siglip (zhai2023sigmoid) encoder for input processing, Phi-3 (abdin2024phi) for inference, and a 2-layer MLP projector (with GELU activation in each layer) for feature mapping and output. The SAVP module utilizes dual-stream input to capture richer deception cues, although it may introduce redundancy. The PVTM mechanism effectively filters out unnecessary tokens, allowing the model to focus on the most important cues, ultimately improving performance.
Detailed Setting of Comparison Methods
In this study, we compare our proposed method with several baseline models, including PatchNet (Wang_2022_CVPR), CoOp (zhou2022learning), IADG (zhou2023instance), LLaVA (liu2023llava), Qwen-VL (Qwen-VL), Minigpt4 (zhu2023minigpt), and Bunny (he2024bunny). All of these comparison methods were implemented using the official open-source code, adhering to the specific settings and configurations provided by the respective authors. For the finetuned MLLMs, the training process was conducted using the dataset we introduced in this work.
Appendix E More Examples and of Experimental Results of FaceShield
To illustrate the effectiveness of the proposed FaceShield model across various tasks, we present example outputs for the coarse-grained classification, fine-grained classification, reasoning, and attack localization tasks in Fig. 18 to 21. These examples highlight the superior performance of FaceShield compared to other models, including LLaVA (liu2023llava), Minigpt4 (zhu2023minigpt), QwenVL (Qwen-VL), and Bunny (he2024bunny).
Coarse-Grained Classification Task (Fig. 18 and 19)
In Fig. 18, FaceShield correctly identifies the face as a spoof by observing critical features such as the unnatural gloss and artificial attributes of the face (e.g., it is a mannequin with painted features). Competing models, such as LLaVA and Qwen, provide inconsistent judgments, while MiniGPT-4 fails to determine the authenticity of the face. Similarly, in Figure 19, FaceShield recognizes the spoof face as being displayed on a screen, correctly leveraging cues such as unnatural lighting and the screen context. Other models, however, misclassify the face as real or fail to reach a conclusion. These results demonstrate that FaceShield excels in identifying subtle spoofing indicators, such as lighting anomalies and material inconsistencies, providing reliable binary classification for real and spoofed faces.
Fine-Grained Classification Task (Fig. 20)
In Fig. 20, FaceShield accurately classifies the spoof type as a “Fakehead Attack,” identifying key features such as lifeless eyes and uniform skin tone, which are typical of a mannequin. Competing models either fail to identify the correct spoof type or provide incomplete and inconsistent descriptions.
Reasoning Task (Fig. 22)
Fig. 22 demonstrate FaceShield’s ability to provide comprehensive and interpretable reasoning for its decisions. In Figure 22, FaceShield uses a structured reasoning framework based on four core attributes: facial lighting, global shape consistency, sense of depth and three-dimensionality, and the presence of edges indicative of spoofing. This structured approach allows FaceShield to conclude that the face is a spoof with high confidence. Competing models either misclassify the face as real or provide reasoning that lacks depth and critical observations.
Attack Localization Task (Fig. 21)
In Fig. 21, FaceShield precisely locates the spoofing region (e.g., the eyes covered by a paper card) and provides accurate bounding box coordinates. In contrast, other models provide vague or incorrect localization results. The precise identification of spoofing areas demonstrates FaceShield’s strength in spatial reasoning and interpretability, which are critical for practical anti-spoofing applications.
Summary of Advantages
Across all tasks, FaceShield consistently outperforms competing models due to its:
  • Accurate Identification: Superior coarse-grained and fine-grained classification of spoofing attacks with fewer errors.
  • Comprehensive Reasoning: Structured and interpretable analysis based on core visual and contextual attributes.
  • Precise Localization: Accurate and detailed identification of spoofing regions, enabling actionable insights.
These examples collectively highlight FaceShield’s capability to address the multifaceted challenges in face anti-spoofing with a high degree of accuracy, interpretability, and robustness.
Furthermore, to highlight FaceShield’s accuracy in classifying on real human face images, we provide additional examples in Fig. 24, further validating the model’s robust capability in handling genuine face data.
Additionally, Fig. 23 presents examples of misclassified cases. Analysis shows that frequent errors occur between Flexible and Rigid Masks, as well as Print and Replay attacks. These issues stem from insufficient data for Flexible and Rigid Masks, making it difficult to map material properties to attack types, while distinguishing Replay and Print attacks requires more contextual information.
Figure 18: Example of Coarse-grained Classification Task
Figure 19: Example of Coarse-grained Classification Task
Figure 20: Example of Fine-grained Classification Task
Figure 21: Example of Attack Localization Task
Figure 22: Example of Reasoning Task
Figure 23: Example of Misclassification
Figure 24: Example of Bonafide Face